OpenStudy (anonymous):

The function y=-0.03x^2+1.60x models the path of a kicked soccer ball. The height is y, the distance is x, and the units are meters. How far does the soccer ball travel? How high does the soccer ball go? Describe a reasonable domain and range for the function.

OpenStudy (anonymous):

well the range can't be negative, because the ball can't go through the ground at any point

OpenStudy (anonymous):

its not going to go straight because gravity exists...

OpenStudy (anonymous):

alright, so what do I do?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

to knw how far the ball has gone, find the x-intercepts

ganeshie8 (ganeshie8):

y=-0.03x^2+1.60x put y = 0 to find x-intercepts 0 = 0.03x^2+1.60x solve x, can u ? :)

OpenStudy (anonymous):

Lol. No. :/

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Heres a hint for \[0=0.03x ^{2}+1.60x\]You can take an x out to get\[x(0.03x+1.60)=0\]Now, can you solve from here?

OpenStudy (anonymous):

No. That just makes it harder

OpenStudy (anonymous):

@ganeshie8

OpenStudy (nincompoop):

http://www.mathsisfun.com/quadratic-equation-solver.html http://www.mathsisfun.com/algebra/quadratic-equation.html http://www.mathsisfun.com/algebra/completing-square.html there's all that you need to solve for x in a quadratic equation no one will be able to help you if you don't start learning some pertinent things

OpenStudy (anonymous):

get the x-intercept as ganeshie8 quoted initially by equation 0=-0.03x^2+1.6x , factor out x to get x(-0.03x+1.6) where either x=0 or -0.03x+1.6=0 and your x-intercept will be x=0 and x=53.33....., get axis symmetry which will be mid-point of the intercepts and that one will give you the height y since for a body in a projectile motion the h-max is half the horizontal range x for reference check projectile motion to confirm that,.

OpenStudy (anonymous):

thankyou. BTW do i do anything after I solve for x?

OpenStudy (anonymous):

get mid-point do you know how to do that?